OpenStudy (anonymous):

Adenan makes a scale drawing of a rectangular flower box. The length is 5 in., and the width is 3 in. Adenan changes the scale of the drawing from 1 in. : 4 ft to 1 in. : 6 ft. Which statement about the dimensions of the flower box is true? The length of the flower box under the new scale is 15 ft. The length of the flower box under the old scale is 30 ft. The width of the flower box under the old scale is 12 ft. The width of the flower box under the new scale is 24 ft.

OpenStudy (solomonzelman):

\(\huge\color{magenta}{ \rm 5 }\) \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\Huge\frac{~~~~~~~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}}\) \(\huge\color{darkviolet}{ \rm 3 }\) lets convert our units all to inches. 4 ft 1 inches = 49 inches 6 ft 1 inches = 73 inches His scale factor changed from 49 inches to 73 inches, that means that all his dimensions are 73/49 times greater...

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

no I didn't get the problem...

OpenStudy (solomonzelman):

according to what I said none of the answers are right....

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

oops meant ik to the i didnt get the answer

OpenStudy (solomonzelman):

It seems as though there isn't a/t wrong with what I said, but apparently, it is wrong.

OpenStudy (anonymous):

oh

OpenStudy (solomonzelman):

sorry

OpenStudy (anonymous):

its k